Grade 10Trigonometry

If sec theta + tan theta = x then what value of sec theta

Rahul rai
7 Years agoGrade 10

1 Answer

Arun
7 Years ago
Dear student
 
As we know
 
sec^2 theta – tan^2 theta = 1
 
hence sec theta – tantheta = 1/x
 
now add both the results
 
2 sec theta = 1/x
 
sec theta = 1/2x